OpenStudy (boxman61):

Divide Complex number 10-5i/3-i (10-5i/3-1)(3+i/3+i) = 30-5i+5/10 = 35-5i/10 Instructors answer is 7/2-(1/2)i need to divide by 5 somehow just not sure how

OpenStudy (anonymous):

your answer is good: 35-5i/10 = 35/10 - 5i/10 = 7/2 - (1/2)i

OpenStudy (boxman61):

so its just reduced by common factor of 5?

OpenStudy (anonymous):

yes...

OpenStudy (boxman61):

thanks